OpenStudy (anonymous):

A distant star has a single planet circling it in a circular orbit of radius 6.33 × 1011 m. The period of the planet’s motion about the star is 828 days. What is the mass of the star? The value of the universal gravitational constant is 6.67259 × 10−11 N · m2/kg2. Answer in units of kg

OpenStudy (badhi):

Since the planet is going on a circular orbit, A centripetal force should exist. This force is generated by the gravitational forces by the two masses. When the circular velocity is \(\omega\), radius is \(r\) The centripetal force \(F_c\) is, \[F_c=mr\omega^2\] Also from Gravitational field theory, \[F_c=\frac{GMm}{r^2}\] Use these equations to find the appropriate answer

OpenStudy (anonymous):

How do I solve for centripetal force if I don't have mass?

OpenStudy (badhi):

Just do some manipulations to the equations to the equations that i've given by combining them. as a hint all I can say is that you do not need to find \(F_c\)

OpenStudy (anonymous):

Oh ok :)

OpenStudy (anonymous):

Wait I have another question since I figure you may know, how do you solve for the period of an orbit?

OpenStudy (badhi):

Period is given. Use it to find the angular velocity - \(\omega\) \(\text{Period}=\frac{1}{\displaystyle\text{frequency}}\) \(2\pi(\text {frequency})=\omega\)

OpenStudy (anonymous):

Well... I got 2.188*^-25 but that seems so wrong ._.